If consumption is C=100+0.75Yd

Taxes is T=50+0.5Y

Export is X=200

Import is M=50+0.25Y

Government spending is G=150

Investment is I=200. Y is domestic income and Y is private disposable income. Determine the surplus or deficit in the government budget at equilibrium


1
Expert's answer
2021-05-13T10:50:02-0400

Given Information

Consumption is C=100+0.75Yd

Taxes is T=50+0.5Y

Export is X=200

Import is M=50+0.25Y

Government spending is G=150

Investment is I=200

Y is Domestic income

Yd is private disposable income


Y = C + I + G + (X - M)

  =(100+0.75Yd) + 200 + (150 - (200 - 50+0.25Y))

  = 300 + 0.75 (Y - T) + 150 -200 -50 - 0.25Y

  = 200 + 0. 75(Y - 50 - 0.5Y) - 0.25y

  = 200 + 0.75Y - 37.5 - 0.375y - 0.25y

Y - 0.125Y = 200

 0.875Y = 200

 Y = 200/0.875

 Y = 228.57

Thus the Equilibrium Domestic income is $228.57 billion


The Goverment budget at equilibrium is

T = 50 + 0.5Y

  = 50 + 0.5 (228.57)

  = 164.285

Thus the goverment collect tax $164.285 billion

Goverment Spending is $150 billion

Govermet surplus = Taxes - Goverment Spending

               = 164.285 - 150 

               = 14.285

Thus the Government is in surplus of $14.285 billion